OpenStudy (gschibby):

\frac{ 1 }{ 3 }*X*A+I _{2}=A ^{T} How do I proceed with this?

OpenStudy (unklerhaukus):

Are those matrices? are you solving for X?

OpenStudy (gschibby):

Yup

OpenStudy (gschibby):

\[\frac{ 1 }{ 3 }*X*A+I _{2}=A ^{T}\]

OpenStudy (unklerhaukus):

what is A?

OpenStudy (gschibby):

\[\left[\begin{matrix}3 & 6\\ 1& 3\end{matrix}\right]\]

OpenStudy (gschibby):

Would it be right to write it like this: \[X=(3A ^{T}-3I _{2})A ^{-1}\]

OpenStudy (unklerhaukus):

yeah

OpenStudy (gschibby):

and then like this: \[X=3A ^{T}A ^{-1}-3I _{2}A ^{-1}\]

OpenStudy (unklerhaukus):

yeah, i think so, you know A so you can work out the right hand side,

OpenStudy (gschibby):

Yup, I'll give it a try ;) Thanks!

OpenStudy (gschibby):

WOHOOO! Got the right answer!^^

OpenStudy (unklerhaukus):

good on ya!